### v2

# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

Abstract
We present a detailed investigation of photometric, spectroscopic, and polarimetric observations of the Type II SN 2023ixf. Earlier studies have provided compelling evidence for a delayed shock breakout from a confined dense circumstellar matter (CSM) enveloping the progenitor star. The temporal evolution of polarization in SN 2023ixf revealed three distinct peaks in polarization evolution at 1.4 d, 6.4 d, and 79.2 d, indicating an asymmetric dense CSM, an aspherical shock front and clumpiness in the low-density extended CSM, and an aspherical inner ejecta/He-core. SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry in the early and late evolution. The argument for an aspherical shock front is further strengthened by the presence of a high-velocity broad absorption feature in the blue wing of the Balmer features in addition to the P-Cygni absorption post 16 d. Hydrodynamical light curve modeling indicated a progenitor mass of 10 with a radius of 470 and explosion energy of , along with 0.06 of , though these properties are not unique due to modeling degeneracies. The modeling also indicated a two-zone CSM: a confined dense CSM extending up to , with a mass-loss rate of 10-2 and an extended CSM spanning from to at least with a mass-loss rate of 10-4 , both assuming a wind-velocity of 10 . The early nebular phase observations display an axisymmetric line profile of [O I], red-ward attenuation of the emission of H post 125 days and flattening in the -band marking the onset of dust formation.
1 Introduction
Massive stars ( 8 , Heger et al., 2003) reach the termination of their evolutionary phase upon exhaustion of their nuclear fuel and undergo core-collapse supernovae (CCSNe). SNe II form a subset of CCSNe that show Balmer features in their spectral sequence (Filippenko, 1997). The progenitors of these SNe are confirmed to be red supergiants (RSGs) through pre-explosion imaging of several SNe II in nearby galaxies (Smartt, 2009; Van Dyk, 2017).
Massive stars are thought to have steady stellar winds driven by radiation pressure (Chiosi & Maeder, 1986). The Galactic RSGs show an observed mass-loss rate in the range of (Beasor & Davies, 2018), due to its dependence on their luminosity and effective temperature. However, early spectral observations of SNe II have shown a short-lived ‘flash’ phase with narrow high-ionization features, indicating a much higher mass loss rate than steady winds (Dessart et al., 2022). Such features were first studied extensively in SN 1998S (Gerardy et al., 2000) as they trace the confined CSM arising from late stages of mass-loss in RSGs (Yaron et al., 2017). A significant fraction of SNe II discovered by Zwicky Transient Facility (ZTF) exhibit such spectroscopic signatures of early interaction during their evolution ( 36%, Bruch et al., 2023).
The effects of a dense CSM driven by high mass loss rates have been known to delay shock breakout (Förster et al., 2018). It occurs when the shock encounters a medium of significantly low optical depth, allowing the radiation to escape in front of the shock (Waxman & Katz, 2017). The breakout is the first electromagnetic signature in CCSNe and happens within a few hours after core-collapse for a typical RSG (Matzner & McKee, 1999). However, if the massive star is engulfed in a compact, dense CSM, the shock breakout leads to an elongated and brighter electromagnetic signature (Ofek et al., 2010; Chevalier & Irwin, 2011), affecting the UV light curves and color evolution (Hiramatsu et al., 2021a; Terreran et al., 2022; Irani et al., 2023). The emergence of high-cadence all-sky surveys (Förster et al., 2018; Subrayan et al., 2023) coupled with prompt follow-up observations has photometrically revealed that fast-rising light curves are primarily reproduced via interaction with CSM. Such an interaction with CSM complicates the classification of SNe II (Singh et al., 2019). SNe II are primarily classified into two main sub-types (Barbon et al., 1979) based on their plateau phase curve decline rates: a Type IIP SN with a slow decline () and a Type IIL SN with a steep linear decline () (Valenti et al., 2016). The primary distinction between Type IIP/L SNe arose from the divide between the outer hydrogen envelope mass (Heger et al., 2003; Eldridge et al., 2018; Hiramatsu et al., 2021b). However, studies by Moriya et al. (2017, 2018); Morozova et al. (2017) have suggested that a higher decline rate during the plateau phase is primarily driven by interaction with CSM, its compactness, and geometry (Andrews et al., 2019; Jacobson-Galán et al., 2022). Chugai (2021) also indicated that higher CSM mass and wind density have been inferred for steeper declining SNe IIL compared to SNe IIP.
Detailed studies of highly interacting SNe II such as SN 2013fs (Yaron et al., 2017), SN 2020pni (Terreran et al., 2022), SN 2020tlf (Jacobson-Galán et al., 2022) etc, have suggested a much higher mass-loss rate for their RSG progenitors, i.e., . Since steady stellar winds cannot propel such a high mass loss rate in an RSG (Puls et al., 2008), the paradigm has shifted towards a more complex scenario in recent years. These mechanisms include eruptive bursts of mass loss (Smith & Owocki, 2006), extensive RSG mass loss through surface instability (Yoon & Cantiello, 2010), wave-driven mass loss during the end stages of nuclear burning (Quataert & Shiode, 2012) and/or turbulent convection driven by dynamical instability (Smith, 2014). The complex mass-loss mechanisms markedly influence the properties of a CCSN through the diverse pathways of forming CSM and its resulting geometrical configuration.
Understanding the explosion geometry of the progenitor of SNe II and its CSM without directly resolving the source is aided by polarimetric observations. Polarimetric studies of CCSNe have indicated that polarization in CCSNe arises from Thomson scattering in the SN atmosphere (Jeffery, 1991; Wang & Wheeler, 2008). Extensive studies by Leonard & Filippenko (2001); Leonard et al. (2006); Nagao et al. (2019) have highlighted numerous SNe II displaying low polarization during the photospheric phase and a subsequent increase in polarization during the transition to the nebular phase, suggesting the presence of an aspherical explosion. Some peculiar SNe II, such as SN 2017gmr (Nagao et al., 2019), have exhibited an early onset of polarization during the photospheric phase, indicating an asymmetric CSM. The interactions in SNe II help probe the geometric configuration of the CSM, as the shock heating of the CSM primarily drives the early-phase luminosity. This allows for probing various configurations of the CSM, including asymmetries resembling a disk, torus, and asymmetric shell, and identifying clumpiness within the CSM (Smith et al., 2015).
The geometry of the CSM helps constrain the physical mechanism of its formation from the progenitor and sheds light on the observed features. Inputs from the line profiles of narrow and intermediate-width features during the early phase also offer insights into the geometry of the CSM, in addition to its density structure (Smith et al., 2011). Imaging polarimetry provides a general overview of the large-scale structure of the explosion, whereas spectropolarimetry aids in investigating polarization across different elemental features to characterize their distribution (Branch & Wheeler, 2017).
This leads us to the nearest SN II in the decade, SN 2023ixf, on which several studies have been published. SN 2023ixf was discovered by Koichi Itagaki (Itagaki, 2023) at JD 2460084.4 and reported on Transient Name Server (TNS) followed by a prompt classification as a Type II SN (Perley & Gal-Yam, 2023) with visible signatures of flash ionization emission features. Early observational studies detailed the evidence of a shock breakout within a dense CSM (Yamanaka et al., 2023; Jacobson-Galan et al., 2023; Smith et al., 2023; Hosseinzadeh et al., 2023; Bostroem et al., 2023; Teja et al., 2023a; Hiramatsu et al., 2023; Zimmerman et al., 2024; Yang et al., 2024; Li et al., 2024). The evidence of asymmetry in the dense CSM was further discussed by (Smith et al., 2023; Vasylyev et al., 2023; Li et al., 2024). The SN was also detected in X-rays 4 days beyond explosion (Grefenstette et al., 2023; Chandra et al., 2023; Zimmerman et al., 2024) as well as in radio (Matthews et al., 2023) along with tight non-detections in mm-wavelengths (Berger et al., 2023) indicating strong signatures of interaction. Pre-explosion imaging studies (Kilpatrick et al., 2023; Pledger & Shara, 2023; Van Dyk et al., 2023; Neustadt et al., 2024; Jencson et al., 2023; Soraisam et al., 2023; Niu et al., 2023; Qin et al., 2023; Ransome et al., 2024; Xiang et al., 2024) revealed the presence of an RSG progenitor embedded in large amounts of dust with indications of periodic variability. However, the inferred progenitor masses vary significantly, ranging from 8 to 24 . Several other studies have established multi-wavelength limits on progenitor activity (Basu et al., 2023; Dong et al., 2023; Matsunaga et al., 2023; Kong, 2023; Panjkov et al., 2023).
Our work presents the multi-wavelength analysis of SN 2023ixf through hydrodynamic light curve modeling until the early nebular phase and long-term spectroscopic and polarimetric monitoring, highlighting the signatures of asymmetries in SN 2023ixf. The data spanning the first 20 days of our study have already been published in Teja et al. (2023a) discussing the early photometric and spectroscopic analysis. We adopt the explosion epoch as estimated by Teja et al. (2023a), i.e., JD 2460083.315, and the epoch described hereafter in the text is described with respect to this date. We provide an overview of the host galaxy’s properties in Section 2, and summarize the details on observational data acquisition and reduction in Section 3. The photometric properties of SN2023ixf are analyzed and discussed in Section 4, followed by a discussion on its spectroscopic evolution in Section 5. Section 6 delves into the temporal evolution of the polarimetric characteristics. Section 7 describes the hydrodynamical light curve modeling to infer the properties of the progenitor and its multi-faceted CSM. We qualitatively discuss the presence of a confined CSM and the signatures of asphericity in the CSM/ejecta in order to understand the progenitor of SN 2023ixf in Section 8 and summarize our results in Section 9.
2 Host Galaxy - M101
SN 2023ixf exploded in the outer spiral arm of the face-on spiral Galaxy, M 101 (NGC 5457), located in the constellation Ursa Major at a redshift of z = 0.0008046 obtained from NASA Extragalactic Database (NED111https://ned.ipac.caltech.edu). SN 2023ixf is the 3rd closest CCSN in this millennium after SN 2004dj (Vinkó et al., 2006) and SN 2008bk (Van Dyk et al., 2012). SN 2023ixf lies proximal to one of the giant H II regions in M 101, namely NGC 5461 (see Figure 1), and in the immediate neighborhood of the H II region #1086 (Hodge et al., 1990). Using emission-line diagnostics of the spectra of H II region #1086, (Van Dyk et al., 2023) estimated an oxygen abundance through of 8.43 12 + log[O/H] 8.86, which equates to a metallicity of 0.10 Z 0.20 close to the site of the SN. We adopt a mean distance of 6.82 0.14 Mpc ( = 29.17 0.04 mag, Tikhonov et al., 2015; Riess et al., 2022), a Galactic reddening of = 0.00770.0002 mag (Schlafly & Finkbeiner, 2011), the host reddening of = 0.031 0.011 mag (Lundquist et al., 2023), and a total reddening of = 0.039 0.011 mag for SN 2023ixf.
3 Data Acquisition and Reduction
3.1 Ultraviolet/Optical/Near-Infrared Photometry
The optical photometric observations of SN 2023ixf were performed using the robotic 0.7-m GROWTH-India telescope (GIT, Kumar et al., 2022) located at the Indian Astronomical Observatory (IAO) in Hanle, India. The observations were carried out in the Sloan Digital Sky Survey (SDSS) filters beginning 2023 May 20 UT. The data were processed with the standard GIT image processing pipeline described in Kumar et al. (2022), and the steps followed are described in Teja et al. (2023a). We also carried out photometric observations in using a 0.36-m Schmidt Cassegrain telescope (Celestron EdgeHD 1400) at the Home observatory in Nayoro, Hokkaido, utilizing a CCD FLI ML1001E camera with an IDAS filter (standard system). The data reduction and aperture photometry were carried out using the software MIRA Pro x64 (Mirametrics, Inc., 2023). The calibration was performed using the stars from APASS catalog (Zacharias et al., 2013). Additionally, optical photometric observations in bands were also carried out using Atik 460 EX Mono CCD mounted on the 0.61-m Vasistha telescope at Ionospheric and Earthquake Research Centre and Optical Observatory (IERCOO), Sitapur, ICSP, Kolkata. Photometric calibrations were done using Tycho software and the ATLAS catalog (Tonry et al., 2018). -band imaging was also carried out using the 0.51-m telescope at Oku Observatory, Okayama, using the SBIG Camera STXL-6303. The data reduction was performed according to the standard procedure using IRAF.
We monitored SN 2023ixf in the near-infrared (NIR) from the Hiroshima Optical and Near-InfraRed Camera (HONIR; Akitaya et al., 2014) mounted on the 1.5-m Kanata Telescope located at Higashi-Hiroshima Observatory, Hiroshima University, Japan. Near-infrared observations were also carried out using kSIRIUS222The design of kSIRIUS is adopted from SIRIUS (the near-infrared simultaneous three-band camera, the near-infrared simultaneous -band imager attached to the Cassegrain focus of the 1.0-m telescope at the Iriki Observatory in Kagoshima, Japan. The NIR data were reduced using standard procedures in IRAF (Image Reduction and Analysis Facility333https://github.com/iraf-community/iraf), and the photometric magnitudes were obtained through the point-spread-function (PSF) photometry using standard IRAF tasks such as DAOPHOT (Stetson, 1987). The photometric calibration was performed using secondary stars from the 2MASS catalog (Skrutskie et al., 2006).
The Ultraviolet Optical Telescope (UVOT; Roming et al., 2005) onboard the Neil Gehrels Swift Observatory (Gehrels et al., 2004) monitored SN 2023ixf extensively beginning May 21, 2023. We reduced the publicly available data obtained from Swift Archives444Swift Archive Download Portal. Photometry was performed using the UVOT data analysis software in HEASoft, following the procedure described in Teja et al. (2022). Upon using a 5 aperture, a significant part of the nearby H II contaminates the flux, especially in the ultraviolet (UV) bands at later epochs ( 20 d). We used the archival data for the host M 101 obtained on Aug 29, 2006, with OBSID 00035892001, available in the Swift archive, as template images for removing the host contribution. The flux obtained at the SN site in the template images is comparable to the late-phase fluxes in all the swift bands. We employed Swift_host_subtraction555https://github.com/gterreran/Swift_host_subtraction code (Brown et al., 2009, 2014a) to remove the host contribution.
3.2 Optical Spectroscopy
Low-resolution optical spectroscopic observations of SN 2023ixf were carried out using the Himalayan Faint Object Spectrograph (HFOSC) instrument mounted on the 2-m Himalayan Chandra Telescope (HCT) at IAO (Prabhu, 2014). The HFOSC observations were performed in grisms Gr#7 (3500-7800 Å, R 500) and Gr#8 (5200-9250 Å, R 800). The spectroscopic data were reduced in a standard manner using the packages and tasks in IRAF with the aid of the Python scripts hosted at RedPipe (Singh, 2021). Optical spectroscopic observations were also performed using the Kyoto Okayama Optical Low-dispersion Spectrograph with optical-fiber Integral Field Unit (KOOLS-IFU, Matsubayashi et al., 2019) mounted at the 3.8-m Seimei Telescope (Kurita et al., 2020) located in Okayama Observatory, Kyoto University, Japan. The KOOLS-IFU observations were carried out using VPH-blue (4100-8900 Å, R 500). The data reduction was performed using the Hydra package in IRAF and a reduction software developed for KOOLS-IFU data666http://www.o.kwasan.kyoto-u.ac.jp/inst/p-kools. Arc lamps of Hg, Ne, and Xe were used for wavelength calibration. We obtained additional optical spectroscopic data using the TriColor CMOS Camera and Spectrograph (TriCCS) installed on the 3.8-m Seimei telescope having a wavelength coverage of 4000–10500 Å (R 700) with grisms g, r and iz. We used the L. A. Cosmic pipeline (van Dokkum, 2001) to remove cosmic rays. We used arc lamps (Hg, Ne, and Xe) for wavelength calibration, and the flux was calibrated using spectrophotometric standards. All the spectra have been continuum calibrated with respect to photometry and corrected for the redshift of the host galaxy.
3.3 Imaging Polarimetry and Spectropolarimetry
We carried out imaging polarimetric observations of SN 2023ixf using HONIR in R-band ( = 0.65 m). The observations comprised a sequence of four position angles of the achromatic Half-Wave Plate (i.e., 0∘, 45∘, 22.5∘ and 67.5∘). The instrumental polarization is about 0.1%, based on the observations of the unpolarized standard star HD 14069, which is consistent with past observations from HONIR (Imazawa et al., 2023). The polarization angle was corrected by observing several polarized standards BD +59°389, BD +64°106 over different nights.
We obtained spectropolarimetric observations of SN 2023ixf over five epochs on 2023 June 14, 15, 16, 18, and 19 using the CCD Imaging/Spectropolarimeter (SPOL; Schmidt et al., 1992) mounted on the 2.3 m Bok Telescope at Steward Observatory, Kitt Peak, Arizona. We used the 600 l/mm grating with a spectral coverage of 4000-7550Å and obtained a spectral resolution of 26.53 Å using a slit size of 0.4 mm (corresponding to 4.1” in the sky). In addition to this, we used a Hoya L38 blocking filter with a cut-off wavelength of 380nm. We used polarized standards (Hiltner 960; HD 155528) to calibrate the polarization angle, with the average discrepancy between the observed and expected polarization angles for multiple standards being less than 1∘ (Schmidt et al., 1992). Furthermore, the unpolarized standard (BD+28°4211) observations confirmed that SPOL exhibited low instrumental polarization, measuring at less than 0.1%. We use the same data acquisition and reduction procedure mentioned in Bilinski et al. (2018). The polarized spectra corresponding to four exposures that sample 16 waveplate positions were flat-field corrected, background subtracted, and flux calibrated using the SPOL polarization data reduction pipeline to obtain the Stokes parameters q, u, p, and polarization angle.
4 Photometric Evolution
Panchromatic light curves covering ultraviolet, optical, and near-infrared wavelengths from various facilities are shown in Figure 2, displaying the well-sampled SN evolution spanning 1.4 d – 150 d after the explosion. All the magnitudes presented in this paper were calibrated to the AB system using the transformations described in Blanton & Roweis (2007).
4.1 Rise times, decline rates and plateau-length
The light curve peaks were measured by fitting the light curves with a smooth spline, calculating the gradient () of the fit, and identifying the zero-crossing points to determine the peaks. SN 2023ixf shows a sharp rise to the peak in the ultraviolet bands - and (i.e., blue peak) with a rise time of 4.5 d. The rise time is 6 d in -band, which is faster than the prototypical rise time of 10 d in Type II SN 1999em (Leonard et al., 2002). The steeper rise to the maximum for SN 2023ixf is attributable to the interaction of the SN ejecta with CSM Förster et al. (2018). The light curves in the redder bands () seem to show a steep rise to a distinct shoulder, i.e., red shoulder post 5 d and a gradual ascent to the maximum, i.e., red peak at 16 d. The red shoulder is seen as an abrupt change of slope in the gradient of the early light curve. The presence of the red shoulder is driven by the flux excess due to the interaction of SN 2023ixf with a confined CSM since it occurs immediately after the peak in the UV wavelengths (i.e., blue peak). A similar shoulder and peak are seen in the simulated red-band light curves of SNe II with strong wind models ( 10-3 ) by Dessart et al. (2017), affirming an alternate scenario to look for signatures of early interaction in the light curves besides the rise time. The rise time for the red peak in SN 2023ixf s higher than that of SNe IIP (7.0 0.3 d) and in the ballpark of SNe IIL (13.3 0.6 d) (Gall et al., 2015).
Post-maximum, the multi-band light curves of SN 2023ixf settle onto a plateau of roughly 75 d in and other redder bands before transitioning to the radioactively-powered tail phase at 90 d. The plateau length is at the shorter end of the typical plateau length of 100 d for SNe II, hence putting SN 2023ixf amongst some of the seldom observed short plateau SNe (Hiramatsu et al., 2021b; Teja et al., 2022). We estimated the plateau decline rates of SN 2023ixf following the prescription of Anderson et al. (2014a). The -band light curve of SN 2023ixf showed an early plateau decline rate (s1) of and a late-plateau decline rate (s2) of . The late-plateau decline rate of SN 2023ixf is higher than the mean decline rate of 1.3 inferred for SNe II (Anderson et al., 2014a). The tail phase decline rate (s3) of SN 2023ixf is which is faster than the characteristic decline rate of 56Co to 56Fe (i.e., 0.98 ) indicating incomplete trapping of and -rays.
4.2 Light curve comparisons with other SNe II
The comparison of -band absolute magnitude light curve of SN 2023ixf is shown in the top panel (A) of Figure 3. Owing to the short plateau length and fast declining nature of SN 2023ixf, we derived a comparison sample consisting of normal Type IIP SNe: SN 2007od (Andrews et al., 2010), SN 2013fs (Bullivant et al., 2018), SN 2016gfy (Singh et al., 2019), 2017eaw (Szalai et al., 2019) and SN 2020tlf (Jacobson-Galán et al., 2022); short-plateau Type IIP SNe: SN 2018gj (Teja et al., 2023b), SN 2020jfo (Teja et al., 2022), and Type IIL SNe: SN 2013by (Valenti et al., 2015), SN 2014G (Terreran et al., 2016), SN 2017ahn (Tartaglia et al., 2021), SN 2018zd (Hiramatsu et al., 2021a) and SN 2020pni (Terreran et al., 2022). The peak -band luminosity of SN 2023ixf is –18.2 mag, brighter than the average peak luminosity of SNe II (i.e., –16.74 1.01 mag) inferred by Anderson et al. (2014a), which is dominated by the population of slow-declining SNe II. The peak luminosity of SN 2023ixf is similar to the SNe IIL, namely SN 2013by, SN 2014G, and SN 2017ahn, but slightly fainter than that of SN 2020pni; however, it is brighter than the majority of normal and short-plateau SNe IIP. The plateau-decline rate of SN 2023ixf is similar to that of SN 2014G ( 1.7 ) and SN 2013by ( 1.5 ). The 75 d plateau length of SN 2023ixf is similar to that of SN 2014G, SN 2013by and SN 2013fs. Overall, SN 2023ixf shows remarkable photometric resemblance in peak-luminosity, plateau decline rates, plateau length, and plateau drop to SN 2013by and SN 2014G.
We further compare SN 2023ixf to SNe II with CSM interaction, followed up extensively in UV by Swift over the last two decades. The data was downloaded from the Swift Optical/Ultraviolet Supernova Archive (SOUSA777https://pbrown801.github.io/SOUSA/, Brown et al., 2014b) and the Vega-mag were transformed to AB-mag for consistency. In the exhaustive sample of SNe II observed by Swift, the and bands exhibit a rapid surge in the flux, brightening by over 3 mags in 3 days, before reaching a peak magnitude of –20 mag at 4.5 d. This prolonged brightening observed in the initial phase of the UV light curve indicates a shock breakout within a compact and dense CSM, leading to a more luminous and elongated shock breakout event (Ofek et al., 2010). Such a distinct signature in the early UV light curves has been observed in only a limited number of SNe II. The detection of a UV burst extending over 1-day in PS1-13arp by Gezari et al. (2015) was the first observation hinting at the possibility of the shock breaking out into a confined CSM. In our comparison sample, only SN 2018zd (Hiramatsu et al., 2021a), SN 2020pni (Terreran et al., 2022), and SN 2020tlf (Jacobson-Galán et al., 2022) showed a similar intensification of the early phase UV light curve.
The evidence of the sharply declining UV flux was seen in the FUV spectra of SN 2023ixf obtained by Teja et al. (2023a) at 23 d, which shows a featureless spectrum owing to the line-blanketing from the iron-group elements (Bufano et al., 2009). Despite this, we still infer a decent contribution ( 10%, see Panel B in Figure 4) of UV flux to the overall bolometric light curve (i.e., a UV excess) during the mid-plateau phase ( 38 d). Our observations indicate continued interaction with CSM despite the absence of discernible signatures in the spectral sequence outlined in Section 5.2. This aligns with the inference from the theoretical models of interacting SNe II by Dessart et al. (2022) that display early interaction frequently display a surplus of UV radiation during late phases.
4.3 Bolometric Light Curve, Temperature and Radius Evolution
We computed the pseudo-bolometric light curves of SN 2023ixf using SuperBol (Nicholl, 2018). The bolometric luminosity, color temperature (), and radius () evolution of the layer of thermalization were estimated using the blackbody fits to the SED of the SN at each epoch. The missing data in certain filters over the intermediate epochs was interpolated using a low-order spline. We computed the bolometric light curve in 3 wavelength bins, i.e., UVOIR (0.16 – 2.35 m), UVO (0.16 – 0.85 m) and OIR (0.38 – 2.35 m) and is shown in Panel A of Figure 4. For the integration, the blue boundary for the UV and optical wavelengths is defined by the blue edge of the Swift band and the band, respectively. The red boundary for the optical and infrared wavelengths is defined by the red edge of the band and the band, respectively. Henceforth, we refer to the UVOIR pseudo-bolometric light curve as the bolometric light curve.
As discussed by Zimmerman et al. (2024), the color temperature and radius evolution shown in Panel C of Figure 4, shows a steep increase in the temperature from around 14,000 K to 35,000 K over a duration of 2.2 0.1 d and a relatively-flat radius evolution at (2.0 0.2) cm (13 1 AU). The extended heating at a near-flat radius, in addition to the prolonged brightening in the UV flux, as discussed in Section 4.2 are signs of further heating upon interaction with the dense CSM. From the photometric observations at 1.1 d, the bolometric luminosity, = (1.28 0.11) , and effective temperature of (14 2) rises ten folds and two folds, respectively, in a span of just 1 d. However, due to the ensued heating, the bolometric light curve of SN 2023ixf flux peaked later on 4.5 d with a luminosity of 2.5 0.3 .
The temporal evolution of UV and NIR fraction of the pseudo-bolometric luminosity is shown in Panel B of Figure 4. The contribution of the UV flux (0.16 – 0.38 m) to the bolometric flux stays at roughly about 85% even until the bolometric maximum (same as the UV peak). If we ignore the early UV data for SN 2023ixf, the OIR bolometric light curve underestimates the bolometric luminosity by an order of magnitude, emphasizing the importance of UV observations of infant CCSNe and its importance to detailed hydrodynamical modeling. During the early nebular phase, the NIR flux contributes almost 60% to the bolometric luminosity whereas the UV flux contributes roughly 5%, highlighting the importance of NIR observations in determining 56-Ni mass in SNe II.
4.4 56Ni Mass
Upon recombination of all the hydrogen in the outer envelope, SNe II post the plateau phase transitions to the tail phase powered by the decay of , which thermalizes the SN ejecta through the emission of -rays and . We computed the 56Ni-mass using the relation postulated by Hamuy (2003). The mean tail luminosity of SN 2023ixf around 145 d yields a mass of . Upon comparison with SN 1987A at a similar phase which synthesized a -mass of 0.075 0.005 (Turatto et al., 1998), a -mass of is estimated for SN 2023ixf. The above techniques assumed complete trapping of -rays by the SN ejecta. However, that is not always true, especially for short-plateau SNe due to a thinner hydrogen envelope (Teja et al., 2022) and may lead to underestimation of the yield. Hence, we modeled the late-phase UVOIR bolometric light curve of SN 2023ixf spanning 100 – 150 d using the analytical formulation from Yuan et al. (2016). We derived a total -mass of 0.059 0.001 and a characteristic -ray trapping timescale of 220 3 d, indicating a short-lived trapping and a shallower envelope. It is important to note that our parameter estimates are based on data spanning up to 150 days, which may result in a slight overestimation of the 56Ni mass. Therefore, we consider this an upper limit, as there are likely additional contributions to the light curve during the tail phase from CSM interaction, as discussed in Section 4.2.
5 Spectroscopic Evolution
5.1 Flash Spectroscopy Phase 8 d
Earlier spectroscopic studies on SN 2023ixf (Yamanaka et al., 2023; Jacobson-Galan et al., 2023; Smith et al., 2023; Bostroem et al., 2023; Teja et al., 2023a; Hiramatsu et al., 2023; Zimmerman et al., 2024; Zhang et al., 2023) showed a prominently blue continuum coupled with highly ionized features of C IV, N IV, C III, N III and He II along with Balmer features beginning 1.1 d from the estimated date of explosion. The features suggest flash-ionization and recombination of the confined dense CSM engulfing the progenitor of SN 2023ixf. Notably, these studies observed an increase in the relative line strengths of the highly ionized species such as N IV and C IV in contrast to N III and C III, respectively, implying a delayed ionization of the CSM. The increase in line fluxes of N IV, C IV and He II noticed in the spectral sequence of SN 2023ixf spanned 4 d from the explosion (Smith et al., 2023), which roughly corresponds to the peak of UV light curves. This delayed ionization is in concurrence with the inference from the increase in temperature evolution in Section 4.3, as it affirms the breakout of the shock from a confined dense CSM. The He II 4686 vanished in the high-cadence spectroscopic observations presented in Bostroem et al. (2023) between 7.6 d to 8.4 d indicating that the flash spectroscopy phase lasted approximately 8 days.
5.2 Photospheric Phase: 8 - 80 d
The complete photospheric phase spectroscopic evolution of SN 2023ixf from 7 d to 82 d is presented in Figure 5. During the transition from the flash phase to the photospheric phase, we see the reminiscence of interaction with the dense CSM in the form of intermediate-width Lorentzian emission along with the emergence of broad P-Cygni features of and (Teja et al., 2023a). This spans until 10 d, beyond which the Lorentzian emission profile vanishes.
Beginning 16 d, we observe a more intricate and complex multi-peaked profile of with a persistent high-velocity absorption feature, which we extensively discuss in Section 5.3. In addition to the distinguishable Balmer lines in the early spectra, several distinct features start to appear 16 d onwards, particularly Ca II H&K and He I 5876 Å. As the photosphere recedes further into the ejecta, we detect the appearance of various lines of Fe II, Sc II and Ba II, and the Ca II Triplet beginning 25 d. These metal features are visibly developed around 36 d. The broad Na ID starts to emerge at the location of He I 5876, evident from the slightly broadened absorption trough at around 40 d (continuing up to 54 d). This is further indicative of the cooling of the ejecta (Gutiérrez et al., 2017a) and is suggestive of an ejecta that is cooling a lot slower than typical SNe II as the appearance of Na ID, is usually observed around 30 d (Gutiérrez et al., 2017a). Towards the end of the plateau drop, i.e. around 70 d, we also begin to observe O I 7774.
In the top-right and bottom-right panels of Figure 5, we compare the spectrum of SN 2023ixf with other SNe II spectra encompassing a sample of SNe IIL and short plateau SNe i.e. SN 2013by (Valenti et al., 2015); SN 2014G (Terreran et al., 2016); SN 2018gj (Teja et al., 2023b); SN 2020jfo (Teja et al., 2022); SN 2020pni (Terreran et al., 2022); SN 2020tlf (Jacobson-Galán et al., 2022). Photospheric features develop pretty late in SN 2023ixf compared to other Type IIL and short-plateau SNe II, significantly later than normal SNe II. In the early plateau phase ( 25 d), the spectra blueward of H is generally dominated by metallic features as in the case for SN 2014G, SN 2020jfo and SN 2020pni; however, the features are relatively underdeveloped in SN 2023ixf and SN 2018gj. The line strengths of absorption features appear much weaker in SN 2023ixf than in spectral lines of other SNe. This plausibly hints at the ejecta being still hot and/or metal-poor compared to normal SNe II. We also observe that SN 2023ixf exhibits a weaker P-Cygni absorption during its early photospheric phase in tandem with short-plateau SNe from (Hiramatsu et al., 2021b) but not with short-plateau SNe 2018gj and 2020jfo. This observation suggests that weaker Balmer absorption doesn’t necessarily result from a lower mass of the hydrogen envelope in short-plateau SNe. The weaker Balmer absorption is also seen in SNe 2014G and 2020pni in our comparison sample. The spectral features are likely less prominent due to the luminosity from interaction enhancing the ejecta’s temperature (and ionization), leading to an ionization wave penetrating the ejecta inwards from the cold, dense shell (CDS, Chevalier & Fransson, 1994). CDS arises from the higher density of the inner CSM (, see Section 7) resulting from the cooling of shocked areas during the early phases. This phenomenon aligns with observations of Type IIL SNe, which typically exhibit weaker P-Cygni absorption troughs since these SNe likely have more significant interaction with CSM than typical Type IIP SNe (Gutiérrez et al., 2014; Polshaw et al., 2016).
5.3 Evolution of H : High-Velocity Absorption and Clumpy CSM
We show the temporal evolution of the and covering the photospheric phase of SN 2023ixf in Figure 6. We observe the emergence of the broad P-Cygni feature of by 7 d in our spectral sequence with a blue edge of 8,500 . A similar broad P Cygni absorption manifests around , confirming the emergence of the SN ejecta. A few epochs later, we observe a high-velocity (HV) broad absorption trough the blue-ward of H P-Cygni absorption beginning 16 d where the two features are merged and cannot be distinguished. This merged broad component slowly migrates to two distinct components in the spectrum on 27 d. The appearance of the HV absorption in H is synonymous with the red peak in light curves of SN 2023ixf. The absorption minima of the HV absorption lies roughly at 13,500 whereas the blue edge of the absorption profile extends up to 20,000 at 16 d. We confirm its association with hydrogen since we see an analogous profile in , although its effect is not as pronounced due to the low optical depth of these lines. This HV feature was also reported by Teja et al. (2023a), who dismissed its association with Si II as it would lead to line velocities lower than the photospheric velocity.
In the case of an expanding SN ejecta, we tend to observe an absorption component forming from the inner layers of the ejecta moving towards our line-of-sight, leading to a P-Cygni profile of H during the plateau phase. The outer recombined ejecta does not contribute towards the absorption profile (Chugai et al., 2007). However, the collision between the SN ejecta and CSM creates a dual-shock structure, where the forward shock moves through the CSM while the reverse shock travels within the SN ejecta (Chevalier, 1982). This results in the ionization of the outer layers of the unshocked ejecta by Lyman- photons, causing the emergence of HV absorption features blueward of the P-Cygni absorption (Chugai et al., 2007). Such an HV feature of hydrogen arising due to interaction is generally narrow and typically starts appearing a month after the explosion (Gutiérrez et al., 2017a). Such HV features are generally narrow and don’t show a considerable evolution in velocity. In contrast, the HV absorption feature in SN 2023ixf is broad (FWHM 7,000 ), starts appearing at 16 d and shows considerable evolution in velocity from 13,500 at 16 d to 9,500 at 70 d. Hence, it is unlikely that the feature arose due to the interaction with the dense CSM. We discuss later in Section 8.4 that the feature likely arises from the freely expanding ejecta across the polar latitudes of the SN, whereas the P-Cygni profile arises from the decelerated ejecta breaking out of the dense CSM.
The P-Cygni profile of H from 10 – 32 d also shows many intricate structures indicating the presence of clumpy matter in the interaction region. We spot many distinct clumpy features at similar velocities in H and H, and they disappear as the SN evolves into the mid-plateau phase ( 40 d), indicating that the SN ejecta overcomes most of the clumps beyond the mid-plateau phase.
The emission peak of has a blueshifted offset by as much as 3,000 at 32 d but evolves towards zero rest velocity by the end of the plateau. This is likely a result of the steep density profile of the ejecta leading to a higher occultation of the receding portion of the ejecta (Anderson et al., 2014b). The blueshifted offset of the emission peak of H (at 30 d) correlates with the decline rate during the plateau phase and the peak luminosity of SNe II (Anderson et al., 2014b). The steep decline rate of SN 2023ixf ( is in agreement with its large offset of 3,000 indicating that the ejecta mass is smaller in SN 2023ixf leading to its shorter plateau. Although this effect is commonly seen in SNe II, this effect is more pronounced in only a handful of events, e.g., SN 2014G (Terreran et al., 2016) and SN 2018gj (Teja et al., 2023b), where the emission peak stays blue-shifted late into the nebular phase.
5.4 Line velocity Evolution
We compare the velocity evolution of Fe II 5169 for SN 2023ixf with other SNe II in Figure 7. We estimated the line velocity evolution from the blue-shifted absorption trough of their line profiles in the redshift-corrected spectral sequence of SN 2023ixf. We adopt the Fe II 5169 velocities as the photospheric velocity since it forms closest to the photosphere (Dessart & Hillier, 2005) and is the least blended among the iron lines seen in our spectral sequence. The photospheric radius estimated from Fe II absorption trough closely mirrors the blackbody radius as shown in Figure 4. During the mid-plateau phase ( 53 d), the Fe II 5169 velocity inferred for SN 2023ixf is 4350 . This positions it at the higher end of the 1- range when compared with the mean velocity of Fe II 5169 for an extensive collection of SNe II by Gutiérrez et al. (2017a, 3537 851 km/s). This is evident in the spectral comparison during the late-plateau phase (see Figure 5) since we observe higher line blending in SN 2023ixf due to its higher photospheric velocity in comparison to all other SNe II in our comparison except that of SN 2014G.
The trend continues onto the early nebular phase ( 116 d), where SN 2023ixf displays a Fe II 5169 velocity of 3330 compared to the mean value of 2451 679 from Gutiérrez et al. (2017a). The higher photospheric velocity is in agreement with the luminosity-velocity correlation of homologously expanding recombination front of hydrogen (Kasen & Woosley, 2009) since SN 2023ixf is brighter during the mid-plateau phase compared to a normal Type II SN (Anderson et al., 2014a). Additionally, the CSM interaction in SN 2023ixf could also drive the elements at outer/faster regions of the ejecta to be reionized and recombined, leading to a higher estimate of photospheric velocity (Andrews et al., 2019).
5.5 Early Nebular Phase 90 d
The nebular-phase spectral sequence of SN 2023ixf is shown in the left panel of Figure 8. The nebular spectra of SN 2023ixf displays prominent emission features of H , Mg I] 4571, H , O I 5577, Na I D / He I, [O I] 6300, 6364, H , [Ca II] 7291, 7324 and the Ca II NIR triplet with a flat continuum, typical of SNe II. We observe signs of asymmetry in the line profile of certain emission features during the nebular phase. We observe a dual-peaked axisymmetric888The separation between the two components is less than that of the two components of the O I doublet profile of the [O I] doublet blueshifted by 1500 in SN 2023ixf during the early nebular phase in Figure 9. In addition, we observe an apparent redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 possibly indicating asymmetries in the ejecta. However, the H and [Ca II] showed a single peak symmetric profile with visible signs of redward attenuation as the SN progressed into the nebular phase.
The emergence of an asymmetric emission profile of H and [Ca II] arising due to the attenuation of the red-ward emission from the receding portions of the ejecta is first noticed in our spectral sequence beginning 125 d in Figure 9. This indicates the onset of dust formation and was first noticed in SN 1987A (Lucy et al., 1989). We also observe that the red-blue asymmetry increases as the SN evolves into the nebular phase, indicating an increased dust formation with time (Bevan et al., 2019). The early signatures of dust suggest its formation is happening inside the CDS since the SN ejecta during the 125 – 140 d is too warm for the condensation for molecules (Kozasa et al., 1991). The flash-ionization features in the early spectral sequence and the steep rise in early UV light curves of SN 2023ixf conclusively indicated the presence of a dense CSM. As the shockwave from the SN encountered the denser CSM, it decelerated, compressing the material and increasing the density within the shocked CSM. Radiative cooling then facilitated the emission of photons, aiding in its cooling and forming a distinctive CDS (Chugai, 2009) in the denser CSM. The CDS enables an additional pathway for dust formation in interacting SNe II (Rho et al., 2018). In addition, the clumpiness within the extended CSM encompassing SN 2023ixf facilitates the formation of additional CDS (in addition to its formation in the dense CSM), consequently enhancing molecule formation and eventually forming dust (Inserra et al., 2011). Since nebular phase H arises from the inner ejecta, it wouldn’t show wavelength-dependent attenuation if the dust is formed in the outer CDS. This suggests that the regions where Balmer lines form and dust formation occurs essentially overlap, indicating thorough mixing of the CDS into the inner ejecta following a significant episode of CSM interaction (Bevan et al., 2019).
We also see evidence of flattening in the -band light curve of SN 2023ixf beyond 125 d in Figure 2 evolving at 1.3 0.1 against the relatively consistent decline of 1.8 0.1 in the and bands. This indicates that the continuum luminosity in NIR is evolving steadily; however, the -band light curve is evolving rather slowly due to the emission from CO overtone around 2.3m and/or formation of warm dust. Although we don’t have spectral conformation in NIR, Type II SN 2017eaw (Rho et al., 2018) showed the presence of the first overtone of CO as early as 124 d and also showed flattening in their -band light curves. This strengthens our inference for indirect conformation of molecular CO formation and eventually dust in the case of SN 2023ixf.
6 Polarimetric Analysis
Figure 10 presents spectropolarimetry observations of SN 2023ixf during the early plateau phase. Due to the low signal-to-noise ratio of the individual epochs, we averaged out the first three epochs (June 14, 15, 16) and the following two epochs (June 18, 19) into two spectropolarimetric observations represented by their average phase of observations. The polarization is expressed as a sum of the and stokes parameters, P , and the polarization angle (PA) is given by the PA = 0.5 .
6.1 Estimating Interstellar Polarization
To analyze the intrinsic polarization of SN 2023ixf, it’s essential to quantify the Interstellar Polarization (ISP) originating from the irregularly shaped dust particles present across the line of sight in both the Milky Way and the host galaxy M 101. To estimate an upper limit on the ISP, we assume a Serkowski-Galactic type ISP (Serkowski et al., 1975) for M 101 by assuming a similar size and dust composition as the Milky Way. Using the extinction values along the line-of-sight of SN 2023ixf in Section 2, the upper limit on the ISP from Serkowski et al. (1975) is 9 = 0.35%. Since the three peaks in our imaging polarization evolution of SN 2023ixf are all higher than the maximum limit on ISP, we reasonably conclude that the observed polarization characteristics originate from the SN itself.
However, to infer any further detail from the polarimetric and spectropolarimetric data, especially during the mid-plateau phase, where the mean continuum polarization seems to drop below the , we must constrain ISP more rigorously from our observed dataset. To extract ISP from our dataset, we assumed that emission peaks of P-Cygni profiles have no intrinsic polarization, similar to several spectropolarimetric studies (Chornock et al., 2006; Nagao et al., 2019). In the case of optically thick lines such as H, multiple scattering processes diminish the geometric characteristics conveyed by the photospheric radiation from the continuum, consequently leading to the depolarization of the emission line (Hoeflich et al., 1996). The emission peak of H and H in Figure 10 exhibits non-zero polarization, featuring an angle distinct from that measured in the continuum range. We estimated the wavelength dependence of the ISP using the averaged spectropolarimetric observations of our entire data set (June 14, 15, 16, 18 and 19) by fitting the wavelength ranges around the peak of H and H with the Serkowski function (Serkowski et al., 1975) i.e. . We fixed the K value to 1.15 (same as Milky Way) and obtained a best-fit for the ISP at and . At the rest-wavelength of 6000Å we estimate an ISP of = -0.07 0.02 %, = 0.22 0.06 % and use this to correct the ISP in our -band imaging polarization evolution and values adopted from Vasylyev et al. (2023) in Figure 11.
6.2 Polarization in the Stokes Q-U Plane and Comparison with other SNe II
In Figure 11, we present ISP-corrected imaging polarization evolution of SN 2023ixf in -band spanning 1.4 d until 140 d. Figure 11 shows 3 distinct peaks in polarization; Point A999We refer to the first observation as point A since we don’t have any data prior to this epoch. [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observed stokes parameters ( and ) in -band are shown in Figure 12. Our first imaging polarimetric observation of SN 2023ixf was conducted at 1.4 d, corresponding to Point A, and is associated with the flash-ionized phase due to the interaction of the SN ejecta with an asymmetric, compact, and dense CSM. This is similar to the value of continuum polarization inferred using spectropolarimetric data by Vasylyev et al. (2023). However, we observe a decrease in the polarization signature observed from the dense CSM on the 2nd epoch, i.e., 2.5 d, which is in stark contrast to the rising luminosity (see a six-fold rise in -band luminosity). However, the continuum polarization reported by Vasylyev et al. (2023) at the same epoch, after correction for ISP, indicates a slight polarisation increase from the first epoch, consistent with rising luminosity driven by interaction. The discrepancy could likely be due to the increased line depolarization of H, leading to a decrease in -band polarization as the PA remains the same.
At the next epoch on 4.4 d, we see a change in PA to 55 which later evolves to the polarization peak B at 6.4 d. The polarization peak B lies towards the end of the flash phase transitioning to the early photospheric phase, which shows interaction signatures in the form of intermediate-width P-Cygni profiles (Smith et al., 2023) and close to the emergence of the broad Balmer features at 7 d (see Section 5.2). Polarization at Peak B likely arises from interaction with extended CSM and/or an aspherical shock structure of SN 2023ixf. Following Peak B, we see a decline in polarization across the early plateau phase ( 22 d); however, the PA continues evolving slowly. There is slight disagreement with the continuum PA reported by Vasylyev et al. (2023) likely due to the depolarization of the emission peak of the P-Cygni profile of H. We do observe a revival in the polarization evolution at 28 d with the PA evolving from 100 to 150 . We have a 40-day gap before our next observations at 70 d, before the end of the plateau phase. The observed polarization evolution of SN 2023ixf after 70 d demonstrated a rise towards Peak C in polarization at 79.2 d towards the end of the plateau phase. Since the innermost layers of the ejecta ( 3000 ) are revealed as a result of the entire hydrogen envelope becoming optically thin, this suggests that the electron-scattering atmosphere of the inner core shows a substantial departure from spherical symmetry. The polarization dropped slowly across the early nebular phase, whereas the PA increased gradually. SN 2023ixf shows two preferred dominant axes: a) along the CSM-outer ejecta, where the SN evolved during the early phase until the epoch of polarization minimum on 21.3 d, and b) along the inner ejecta/core, where the SN evolved during the late-plateau and early nebular phase, which is inferred from its polarization evolution.
In a recent study, Nagao et al. (2024) compared the intrinsic polarization properties of 15 SNe II. Broadly, it was inferred that the degree of polarization rises significantly () in the transition/nebular phase; however, it is usually lower during the photospheric phase (). Anomalous cases exist in SN 1987A (Jeffery, 1989) and SN 2021yja (Vasylyev et al., 2024), where a significant polarization is observed during the early plateau phase, gradually declining towards the end of the plateau phase. This is likely driven by electron scattering from interaction with an aspherical CSM, which is analogous to the early continuum polarization (Point A) in SN 2023ixf. However, we also see a secondary peak (Peak B) in polarization in SN 2023ixf, which could possibly arise from interaction with a clumpy low-density extended CSM and/or an aspherical shock structure (see Section 8.4 for further discussions). Towards the end of the plateau phase, the swift increase in polarization during the late-plateau phase and steady decrease during the early nebular phase is similar to that seen in several SNe II like SN 1987A (Jeffery, 1989), SN 2004dj (Leonard et al., 2006) and SN 2013ej (Kumar et al., 2016). Such a polarization evolution is often interpreted as arising from the asymmetries in the inner He-core, and the trend seen is similar in SN 2023ixf. Type IIL SN 2017ahn, which showed early flash-ionization features similar to SN 2023ixf, showed low continuum polarization during the late-plateau phase photospheric and nebular phases, indicating that it might have been asymmetric; however, its viewing angle was lying across its polar axis (Nagao et al., 2021). However, it is clear from the observations of SN 2023ixf that the inclination differs from the axis of symmetry.
7 Light Curve modeling
To estimate the explosion parameters and infer the properties of CSM around SN 2023ixf, we performed numerical light curve modeling of SN 2023ixf. We used one-dimensional multi-frequency radiation hydrodynamics code STELLA (Blinnikov et al., 1998, 2000, 2006) for this purpose. Moriya et al. (2023) computed light-curve models of SNe II with various progenitor masses, explosion energies, and CSM density profile and its extent using STELLA. The model grid adopts the solar-metallicity RSG progenitor models with the ZAMS masses of 10, 12, 14, 16, 18 from Sukhbold et al. (2016) and covers explosion energies from to . The CSM is constructed by assuming the mass-loss rates of with a terminal wind velocity of . The density structure of the wind above the progenitors is characterized by the parameter ranging from to . shapes the velocity/density profile of the wind and is inversely proportional to wind acceleration (Moriya et al., 2017). We refer the reader to Moriya et al. (2023) for additional details regarding the model grid. The extensive details regarding the model comparison will be presented in Moriya & Singh (2024).
Using the pre-computed model grid, we first looked for the best matching models in the g and r bands by searching for models with the minimum . We found that the models from a low-mass progenitor () best fit the light curves of SN 2023ixf. The best fitting models have the explosion energy of around , the mass of around , the mass-loss rate of around 10-2 , the CSM radius of around , and of around . We performed additional numerical light curve calculations to fit the early UV luminosity and found that the model with the CSM radius of matches the observed dataset, including the photospheric velocity (shown in dashed lines in Fig. 13). The mass of the confined CSM below in this model is . This extra model assumes that is uniformly mixed in the entire ejecta. We found that the fully mixed model better fits the transition from the plateau phase to the tail phase. However, they also lead to an excess emission in the redder bands during the late plateau phase as starts diffusing out earlier than in the case of a centrally concentrated 56Ni. A much more prominent effect of -mixing was seen in the case of SN 2009ib (Takáts et al., 2015) and SN 2016gfy (Singh et al., 2019). Since the 56Ni synthesized in SN 2023ixf is high in comparison to normal SNe II (0.03 , Anderson et al., 2014a), the late-plateau luminosity bump is much more prominent in the redder bands of our models.
We observed a notable decline in the UV flux around 40 days in the best-fitting model (having only an inner dense CSM) presented above in Fig.13. To address this decrease and account for the sustained UV excess in the late phase of SN 2023ixf, we modeled the light curves with an extended CSM component situated right outside the inner CSM from to . This additional CSM component possesses a mass of 0.025 and has a density structure proportional to . The low-density extended CSM component helps sustain a prolonged UV brightness until the early nebular phase. Assuming a constant wind velocity of , the outer CSM component is shaped by an average mass-loss rate of 10-4 . The mass-loss rate should increase towards the time of the explosion to form a density structure proportional to . We found that the shallower CSM structure, such as that proportional to , gives a UV flux excess much higher than observed. Using the observed wind velocities of 50 – 115 (Jacobson-Galan et al., 2023; Smith et al., 2023; Zhang et al., 2023) instead of an assumed value, we infer a mass-loss rate of for the inner dense CSM and for the extended CSM.
SN 2023ixf has a higher peak luminosity, higher 56Ni-mass, and a higher photospheric velocity than a typical Type II SN. The peak -band luminosity of SN 2023ixf is -18.2 mag, which is 4 times brighter than compared to the luminosity of a typical Type II SN, i.e., -16.7 mag (Anderson et al., 2014a). The 56Ni-mass of SN 2023ixf is 0.059 0.001 which is 80% higher than the 56Ni-mass estimate of a typical Type II SN, i.e., 0.033 from (Anderson, 2019). Previous studies on SNe II (Hamuy, 2003; Pejcha & Prieto, 2015; Gutiérrez et al., 2017b) have shown that more energetic explosions lead to higher photospheric velocities and a higher 56Ni-yield in the SN. The estimated 56Ni-mass does indicate that SN 2023ixf is a highly energetic event and is in concordance with the high explosion energy of 2 10 estimated from the light curve modeling. Other works on numerical modeling of SN 2023ixf also suggest high explosion energies. Bersten et al. (2024) estimate 1.2 foe as explosion energy for a 12 RSG with 10.9 as the final progenitor mass. Hiramatsu et al. (2023) estimated 1 foe as explosion energy while exploring 2 foe as well for a 12 RSG with 11 as the final progenitor mass.
In our case, where we have the best match with a lower mass progenitor, slightly higher explosion energies must be required to match the observed light curves. Although we fit both photospheric velocities and light curves simultaneously, the mass and radius of the progenitors are fixed since we adopt progenitor models from Sukhbold et al. (2016). Given the inherent degeneracies between the ejecta mass, radius, and explosion energy (Goldberg et al., 2019), our inferred properties likely represent one possible set of solutions rather than a unique one. However, the high-explosion energy cannot entirely explain the peak luminosity of SN 2023ixf. It is enhanced further by early interaction with the confined CSM. The photospheric velocity of SN 2023ixf is 20% faster than a proto-typical Type II SN at 50 d (see Section 5.4), and a power-law fit to photospheric velocity evolution returned exponent -0.47 0.04, which is slower than the average value derived from a large sample of SNe IIP (–0.581 0.034) in Faran et al. (2014). However, since SN 2023ixf is a short-plateau SN, the slightly low-ejecta mass could be one reason for its higher photospheric velocity (Teja et al., 2022).
8 Discussion
8.1 Timeline of significant epochs for all historic observations
Measurement | Phase (or Phase Range ) | JD | Reference |
Epoch of Explosion | – | 2460083.315 | 1 |
Epoch of Discovery | 0.9 d | 2460084.227 | 7 |
Epoch of Polarization Point A | 1.4 d | 4, This paper | |
Epoch of Shock Breakout | 2.2 0.2 d | This paper | |
2.75 0.3 d | 3 | ||
Epoch of X-ray Detection (NuSTAR) | 3.9 d | 6 | |
Epoch of X-ray Detection (Swift-XRT) | 4.25 d | 3, 10 | |
Change in the profile of the Narrow feature | 4.4 d | 2 | |
Peak in UV Luminosity (blue peak) | 4.5 0.5 d | 3, This paper | |
Peak in He II line flux | 4.5 0.5 d | 3 | |
Transition in PA of Polarization | 4.4 – 4.6 d (starts at 3.5 d) | 4, This paper | |
Epoch of Polarization Peak B | 6.4 d | This paper | |
Emergence of Broad absorption | 7 d | 1, 2, 5, 8 | |
End of Flash-Ionisation Phase | 8 d | 1, 5, 8 | |
Peak of Soft X-ray Emission | 10 d | 3, 9, 10 | |
Disappearance of Intermediate-Width Lorentzian Features | 10 d | 1 | |
Peak in Luminosity of redder bands (red peak) | 16 d | This paper | |
Emergence of Broad HV absorption | 16 d | This paper | |
Epoch of Millimeter Non-Detections (230 GHz) | 2.6 – 18.6 d | 11 | |
Epoch of Radio Detection (10 GHz) | 29.2 d | 12 |
(1) Teja et al. (2023a) (2) Smith et al. (2023); (3) Zimmerman et al. (2024); (4) Vasylyev et al. (2023); (5) Jacobson-Galan et al. (2023); (6) Grefenstette et al. (2023); (7) Itagaki (2023); (8) Bostroem et al. (2023); (9) Chandra et al. (2023); (10) Panjkov et al. (2023); (11) Berger et al. (2023); (12) Matthews et al. (2023)
In its infancy, SN 2023ixf showed a rapid evolution spearheaded by the appearance of an increase in ionization in the early flash-ionization phase, accompanied by a rapid ascent in the early UV flux. Additionally, the increase in the color temperature to approximately 35,000 K until around 2.2 days (see Figure 4) signifies that the shock breakout was delayed and occurred within a confined, dense CSM. Our estimate for the epoch of shock breakout differs slightly from that of (Zimmerman et al., 2024). This variation arises because they use the blackbody radius and the shock velocity approximated from the photospheric velocity to compute the extent of the dense CSM. The epoch of peak UV luminosity (and the bolometric peak) at 4.5 d is synonymous with the peak of the He II line flux of Zimmerman et al. (2024), which traces the strength/flux of the flash-ionization features. The change in the line profile of the narrow feature and the drop in its strength seen in the high-resolution spectroscopy of SN 2023ixf (Smith et al., 2023) also happens around 4.4 d, and is synonymous with UV peak. This indicates that the thermal heating and ionization continued beyond the shock breakout (2.2 d). The emergence of a CDS (Chugai, 2009) within the post-shock CSM and the decelerated SN ejecta is probably contributing to the photo-ionization and prolonged heating observed in the flash-ionized features of SN 2023ixf. During the breakout phase, the radius of thermal emission remained relatively constant at (2.0 0.2) cm (or 13 1 AU), indicating the location where thermal radiation originates from within the dense CSM (Chevalier & Irwin, 2011). This radius, derived from blackbody fits to UV-Optical-NIR data, is typically smaller (as it generally forms deeper) than the radius of photospheric emission () (Moriya et al., 2011) and the surface of last scattering of the dense CSM.
The first detection of SN 2023ixf in X-rays from NuSTAR on 4 d showed a large column density of absorption, consistent with arising from a shocked dense CSM (Grefenstette et al., 2023). However, the next epoch of X-ray observations at 11 d and 13 d (Grefenstette et al., 2023; Chandra et al., 2023) exhibited a substantial decline in the column density of absorption. SN 2023ixf showed the emergence of broad P-Cygni of H at 7 d, and the end of the flash-ionized phase lasted 8 d. We see the appearance of the broad HV absorption of H at 16 d in synonymity with the red peak in our multi-band light curves in Section 4.1. The intermediate-width Lorentzian features from CSM interaction disappeared in the spectra around 16 – 18 d (Smith et al., 2023). SN 2023ixf was not detected at millimeter wavelengths from 2.6 – 18.6 d (Berger et al., 2023). SN 2023ixf was eventually detected in radio wavelengths rather feebly after 29.2 d (Matthews et al., 2023). The timeline of significant epochs for SN 2023ixf is tabulated in Table 1 for reference.
8.2 Progenitor of SN 2023ixf
Numerous works on SN 2023ixf have estimated the progenitor mass, mass-loss rate, and extent of the CSM encompassing the progenitor. Pre-explosion imaging at the site of SN 2023ixf through HST and Spitzer revealed a point source consistent with an RSG enshrouded by a large amount of dust (Soraisam et al., 2023; Jencson et al., 2023; Neustadt et al., 2024). However, no counterpart was discovered in UV or X-rays (Basu et al., 2023; Matsunaga et al., 2023; Kong, 2023; Panjkov et al., 2023). There is, however, a disparity in the estimates of progenitor mass from the pre-explosion imaging revealing estimates in 2 broad ranges, i.e. (9 – 14 , Kilpatrick et al., 2023; Pledger & Shara, 2023; Van Dyk et al., 2023; Neustadt et al., 2024, and (17 – 22 , Jencson et al., 2023; Soraisam et al., 2023; Niu et al., 2023; Qin et al., 2023. Our numerical hydrodynamical modeling best matched a ZAMS progenitor mass of 10 for SN 2023ixf having a radius of 470 . Only other work that performed hydrodynamical modeling of the complete light curve until the early nebular phase indicated a 12 progenitor (Bersten et al., 2024). The short-plateau nature of SN 2023ixf indicates a relatively lower ejecta mass, which is also reflected in its steep plateau decline rate. Furthermore, the considerable blueshifted offset observed in H ( 3,000 ) during the early phase reinforces this deduction, indicating an escalated degree of stripping undergone by the 10 progenitor of SN 2023ixf. A summary of the various estimates of the progenitor mass and the extent and the mass-loss rate of the dense CSM is shown in Figure 15.
Pre-explosion observations of the progenitor of SN 2023ixf revealed variability in the mid-IR and near-IR observations from Spitzer and ground-based telescopes (Kilpatrick et al., 2023; Soraisam et al., 2023; Jencson et al., 2023). However, despite this variability, there is no indication of pre-SN outbursts in the pre-explosion imaging Jencson et al. (2023); Hiramatsu et al. (2023); Ransome et al. (2024), nor any signs of variability in the optical spectrum (Dong et al., 2023; Ransome et al., 2024; Neustadt et al., 2024), which likely denies the existence of episodic mass loss in SN 2023ixf. This lack of pre-explosion outbursts, coupled with the presence of an asymmetric confined CSM (see section 8.4), suggests that the progenitor of SN 2023ixf likely had a rapid rotation and/or underwent a pre-SN interaction with a binary companion (Smith, 2014; Matsuoka & Sawada, 2023). Such an interaction would have led to an enhanced mass loss in the lead-up to the explosion and drive a significant asymmetry in the observed CSM, as inferred from our polarization observations. SN 2023ixf is thus a low-mass RSG progenitor showcasing a multi-faceted CSM geometry arising from enhanced mass loss during its twilight years. The presence of the two-zone CSM in SN 2023ixf (see Section 13) supports the notion of an increased mass-loss rate of the RSG progenitor of SN 2023ixf as it evolved towards its explosion.
8.3 Characteristics of circumstellar material around SN 2023ixf
SN 2023ixf showed several signs of an interaction with CSM both photometrically and spectroscopically. The hydrodynamical modeling in Section 7 emphasized the presence of a confined dense CSM responsible for the origin of the flash-ionization features, steep rise in UV flux, bolometric luminosity and temperature, and an extended low-density CSM, which led to the late-phase UV excess and the clumpy features around H. This brings forward the argument that the progenitor of SN 2023ixf had an enhanced wind that developed shortly before the explosion, leading to the delayed shock breakout.
The modeling also revealed that the true extent of the confined CSM is roughly 5 cm (33 AU), more significant than the nearly-flat radius of thermal emission during the early SN evolution in Section 4.3. Our estimates align well with the estimates from comparison with CMFGEN models (Jacobson-Galan et al., 2023), high-cadence early spectroscopy (Bostroem et al., 2023), early light curve modeling (Hiramatsu et al., 2023) and pre-discovery photometry close to the explosion (Li et al., 2024). Assuming a wind-velocity of 10 , the confined CSM is characterized by a wind-like structure with a mass-loss rate of 10-2 which is consistent with estimates derived using comparison with CMFGEN models (Jacobson-Galan et al., 2023) and narrow emission features (Zimmerman et al., 2024). The mass-loss rate indicates a progenitor star in the eruptive phase, which could lie anywhere between 10-2 – 0.1 (Smith, 2017). The absence of SN 2023ixf detections in mm-wavelengths by Berger et al. (2023) within the first 4 days, suggests a mass-loss rate of 10-2 , aligns with our modeling estimates and is likely influenced by a strong dominance of free-free absorption in the dense CSM along our line-of-sight.
The first detection of SN 2023ixf in X-rays from NuSTAR on 4 d displayed a large column density of absorption, consistent with arising from a shocked confined dense CSM (Grefenstette et al., 2023). However, the next epoch of X-ray observations at 11 d and 13 d (Grefenstette et al., 2023; Chandra et al., 2023) exhibited a substantial decline in the column density of absorption arising from 3 10-4 (Grefenstette et al., 2023). The disparity between mass-loss rate estimates from X-ray observations and other measurements is likely due to the X-rays probing the interaction with the low-density extended CSM, whereas the others are probing the confined CSM. The lower mass-loss rates inferred from X-ray observations are consistent with our findings of interaction with a low-density extended CSM and our mass-loss rate estimate of 10-4 ) from light curve modeling. The non-detection of SN 2023ixf in mm-wavelengths by Berger et al. (2023) after the first 4 days hints towards a weak synchrotron emission from the low-density extended CSM (Hiramatsu et al., 2023).
8.4 Asphericity in the CSM / Outer Ejecta: Asymmetric confined dense CSM and Aspherical Shock front/extended CSM
The temporal evolution of polarization from close to the explosion (1.4 d) up until the middle of the plateau phase suggests significant departures from global sphericity for SN 2023ixf, which encompasses the confined CSM, outer ejecta and quite likely the extended CSM.
8.4.1 Early polarization evolution: Point A
Our initial polarization observation (Point A; 1.1%) at 1.4 d post-explosion aligns with the appearance of flash-ionization lines in the spectral sequence of SN 2023ixf, a consequence of its interaction with the confined dense CSM. Given that the interaction with the dense CSM predominantly influences the luminosity during this phase, the inferred polarization during this period reflects the aspherical geometry of the dense CSM at a PA of 150 ∘. Early epoch polarization values are consistent with the findings of Maund et al. (2023) and Vasylyev et al. (2023). Since the shock breakout doesn’t happen until 2.2 d, the shock-powered SN ejecta lies within the compact dense CSM at 1.4 d and indicates an asymmetric and confined dense CSM surrounding the progenitor of SN 2023ixf. During this phase, the narrow emission of , arising from the pre-shock CSM, shows a blueshifted and asymmetric feature (Smith et al., 2023), potentially caused by the plane of the dense CSM aligning across our line of sight, obscuring the redshifted portion of the CSM.
Owing to the asphericity of the confined dense CSM and the arguments above, we consider the similarity of the CSM around SN 2023ixf to PTF11iqb (Smith et al., 2015) and SN 1998S (Gerardy et al., 2000), which displayed an equatorial geometry of their CSM (i.e. a disk/torus). Similar to PTF11iqb and SN 1998S, SN 2023ixf exhibited a bright UV peak (see Section 4.2), flash-ionization features (see Section 5.1) and show Point A in polarization (PA 150∘, see Section 6) within the first four days. However, this phase in SN 2023ixf in similar in comparison to SN 1998S ( 8 d) and longer than that of PTF11iqb ( 3 d) (Jacobson-Galán et al., 2024). This likely indicates the confined CSM in SN 2023ixf has a large radial extent, i.e., 30 AU, Section 7). Assuming an oblate geometry for the CSM, the polarization at Point A corresponds to a major-to-minor axis ratio of 1.2 (Hoflich, 1991).
8.4.2 Photospheric phase polarization evolution: Peak B
We observe a shift in the imaging polarization evolution at 4.4 d, suggesting a transition in the primary source of continuum luminosity, which was earlier dominated by interaction with the dense CSM. The aforementioned epoch aligns with the peak ionization and UV flux (refer to Table 1 for the timeline); this PA change indicates an alternate source of polarized light. Moving forward, we observe a polarization rise to another point at 6.4 d (Peak B; 0.54%). The polarization is likely a result of the recombination emission from an aspherical hydrogen envelope or additional luminosity from interaction with an asymmetric CSM. Following Peak B, we see a steady decline in polarization while the PA continued evolving slowly across the early plateau phase, falling to as low as 0.11% on 22 d. The epoch of polarization minima coincides with the changing slope during the plateau phase (). Our next epoch of imaging polarization at 28 d sandwiched by two epochs of spectropolarimetric observations (at 27.3 d and 30.9 d), which showed an increase in polarization with the PA changing 50∘ from the polarization minima at 22 d. The early evolution of SN 2023ixf until the middle of the plateau phase showcases the dominant axis in the CSM-outer ejecta, which is illustrated in the q-u plane in Figure 12. Such a preferred alignment onto a dominant axis is a marker of global axisymmetry (Tanaka et al., 2017). Since Point A and Peak B arise from a roughly axisymmetric geometry of the CSM and outer ejecta, the SN is viewed from an angle of inclination of 60 degrees to the plane of the sky. Larger viewing angles will require more significant asymmetries since the polarization will scale as , where is the angle measured from the pole (Hoflich, 1991; Wang et al., 2001). We explore the origin of Peak B from two potential scenarios outlined below:
Interaction with an extended CSM which is asymmetric: There is a clear indication that SN 2023ixf also harbors a low-density extended CSM due to visible signatures of intermediate-width features until 10 d. In addition, we observe clumpiness in the extended CSM through the line profile from 10 – 32 d, which could cause polarized emission. It is also reflected in the requirement for an extended CSM in the hydrodynamical modeling of the UV excess spanning up to the early nebular phase. However, the above effects are likely not strong enough to account for Peak B polarization unless the extended CSM exists as dense clumps in an effervescent zone as prescribed by Soker (2023). In addition, (Smith et al., 2023) noted that the narrow emission from migrates towards a narrower symmetric feature past 4.4 d, indicating that the pre-shock CSM it is likely originating from, is symmetric considering its line geometry. We also observe the effects of UV excess from the interaction with the extended CSM interaction span until the nebular phase; however, the polarization drops drastically to a minimum at 22 d. Additionally, the gradual evolution of the PA during the photospheric phase further weakens the argument that interaction with the extended CSM is the sole cause of this polarized light.
Asphericity in the hydrogen envelope: We see the emergence of the broad P-Cygni in SN 2023ixf at 7 d, which coincides closely with the epoch of Peak B in polarization. This likely indicates that the polarization might arise due to the asphericity in the hydrogen envelope of the SN ejecta since the photosphere would reside in the hydrogen envelope during the photospheric phase. Additionally, we observe the appearance of a broad high-velocity absorption feature blueward of H and H, alongside an absorption feature associated with their P-Cygni profiles. The PA change of 90 degrees from Point A to Peak B emphasizes that the faster-moving ejecta (Peak B) lies perpendicular to the equatorial CSM (Point A). The switch in PA was also promptly observed between 3.5 and 4.6 d in the spectropolarimetric observations of (Vasylyev et al., 2023) wherein the PA of H at 3.5 d becomes the PA of the continuum at 4.6 d. This likely means that such a transition occurred earlier for Balmer line emission than for the continuum along the CSM-outer ejecta dominant axis. We hence firmly believe that the source of the dual broad Balmer absorption features and polarization Peak B arises from the aspherical shock front in SN 2023ixf, induced by its interaction with the equatorial dense CSM. We explain this further in the following subsection.
8.4.3 Aspherical shock structure in SN 2023ixf
Consider the case of an RSG undergoing an explosion surrounded by a torus-shaped dense CSM across the equatorial plane along with a clumpy low-density extended CSM as shown in Figure 14. The shock traversing across the equatorial plane would be decelerated by the confined dense CSM, whereas the shock moving along the higher latitudes will expand freely owing to a low-density CSM, causing an aspherical shock front. Hence, such a geometry would likely allow for two broad absorption features in the spectra of SN 2023ixf arising from the decelerated ejecta and the freely expanding SN ejecta (see Figure 6). The decelerated ejecta broke out from the dense CSM along the equatorial latitudes at 2.2 d. However, the unimpeded, freely expanding ejecta along the polar latitudes required it to traverse the lateral extent of the decelerated ejecta to become discernable later in the observations of SN 2023ixf. It is reasonable to anticipate a counterpart in the emission from the freely expanding ejecta. The distinct visibility of the HV absorption in our spectra implies a component of the freely expanding ejecta moving toward our line of sight. Consequently, this orientation would mask the redward emission, as it would be hindered by its motion away from our line of sight and potentially be obscured by the equatorial decelerated ejecta. When the broad HV absorption feature emerges, we observe a faint signature of extended red edge of the H emission feature around 16 d (see Figure 6).
An aspherical shock structure is crucial for prolonging the survival time of dust grains, resulting in time-varying extinction as was presented in the pre-discovery photometry of SN 2023ixf, 1 day from the explosion (Li et al., 2024). In addition, the aspherical shock front is consistent with a prolonged and weak initial breakout flash in SN 2023ixf and a rising luminosity produced as ejecta flung out after breakout, expanding, and cooling. The asymmetric nature of the dense CSM also explains the low X-ray luminosity of SN 2023ixf in comparison to interacting SNe II (Panjkov et al., 2023), in tandem with the highly asymmetric SN 2017hcc, which exhibited a bright optical-IR emission but a weak X-ray emission (Chandra et al., 2022). Panjkov et al. (2023) described that the stacked X-ray Swift-XRT flux arises from a two-component bremsstrahlung emission consisting of a heavily absorbed hotter component and a less-obscured cooler component, which fits the description of an aspherical shock front in SN 2023ixf.
Grefenstette et al. (2023) and Chandra et al. (2023) highlighted the discrepancy in shock breakout velocities derived from X-ray observations in contrast to the estimates based on optical observations by Jacobson-Galan et al. (2023) and Zimmerman et al. (2024). X-rays in CCSNe are predominantly produced through shock interaction with nearby CSM, and the geometry of the CSM plays a significant role in determining the probability of X-rays escaping and their detectability in an SN. Specifically, an asymmetric CSM structure, like a disk or toroid, facilitates the escape of X-rays along higher latitudes, in contrast with a spherical CSM, where their escape would be impeded. In addition, the mass-loss rate derived by Grefenstette et al. (2023) implies lower optical depth ( 1) to electron-scattering in contrast to the presence of Lorentzian features in the optical spectral sequence until 10 d. The asymmetry in the dense CSM and the aspherical shock structure interacting with the extended CSM likely drove the origin of the X-rays from a region distinct from the UV-Optical-NIR emission. They manifested as lower shock velocity measured by Grefenstette et al. (2023); Chandra et al. (2023) using X-rays compared to the H velocity from our spectral sequence.
The primary source of optical emission is driven by the shock decelerated by the equatorial CSM ( 8,500 ), forming the primary P-Cygni Balmer profile. In contrast, the X-ray observations by NuSTAR and Chandra at epoch 11 – 13 d likely originate from the faster shock ( 13,500 ) arising from the freely expanding ejecta interacting with the low-density extended CSM, also leading to the HV Balmer absorption. A weak-synchrotron emission arising from interaction with the low-density extended CSM (Hiramatsu et al., 2023) is consistent with the lack of detections in mm-wavelengths by (Berger et al., 2023) post 4 days. The mass-loss rate estimates from X-ray observations are consistent with our estimate for the extended CSM (see Section 7. However, Chandra et al. (2023) estimated that the radial evolution of X-ray emission past 13 d evolved as is consistent with a wind profile of whereas our hydrodynamical modeling estimated a steeper wind profile of for the extended CSM, indicating a slight disagreement (see Section 7). However, it is important to acknowledge that neither the 1-D hydrodynamical modeling nor a single component X-ray emission fully captures the complexities of asymmetry in the CSM/ejecta of SN 2023ixf as observed through its polarization evolution. Initial attempts towards a two-component X-ray emission from (Panjkov et al., 2023) attribute it to the asphericity in SN 2023ixf, potentially explaining its underluminous nature. Future studies with 2-D/3-D hydrodynamical modeling with the wealth of information on SN 2023ixf will be necessary to address these details and further our understanding of its progenitor system.
8.5 Geometry of the Inner Ejecta / He-Core: Peak C
Post the spectropolarimetric observations at 31 d, we see an increased polarization in the next epoch of imaging polarimetric observation at 68 d (0.44 %), which lies closer to the end of the plateau phase. There’s a slight rise in polarization towards the 3rd peak (Peak C; 0.5%) at 79.2 d, which lies during the transition from the plateau phase. As SN 2023ixf evolved into the nebular phase, the polarization showed a gradual decline in evolution. This decline is likely attributed to the expansion of the inner He-core, resulting in a reduction of its optical depth to electron scattering (Leonard et al., 2006), consequently leading to a decrease in polarization. However, we also see an evolution of the PA past Peak C from 18 ∘ to 106 ∘, indicating that the scenario is more complex. The evolution of SN 2023ixf past the middle of the plateau phase until the early nebular phase aligns across alternate dominant axes showing axisymmetric inner ejecta / He-core (see Figure 12).
We observe signs of asymmetry in the early nebular phase spectroscopy in the form of redshifted excess in H and [Ca II] 7291, 7324 at around +5,000 in Figure 9. Additionally, a dual-peaked axisymmetric profile of the [O I] doublet, blueshifted by 1500 , is also observed in SN 2023ixf during the early nebular phase in Figure 9. It is important to note that these signatures of asymmetry are distinct from the blueshifted excess seen close to the emission peak of the line profiles of H and [Ca II] 7291, 7324 which is likely a result of dust formation (see Figure 9). In the classification of [O I] line profiles in CCSNe, as proposed by Taubenberger et al. (2009), the double-peaked structure of the [O I] lines exhibiting axisymmetric blueshifted and redshifted peaks is associated with a toroidal distribution of oxygen. Such excitation of oxygen is expected from an asymmetric distribution of 56Ni (Gerardy et al., 2000) when observed from approximately 60 to 90 degrees with respect to the symmetry axis (Maeda et al., 2008; Fang et al., 2024). The line profile of [O I] doublet in the early nebular phase conforms to that picture, echoing the polarization signature detected during the early nebular phase. The effect of asymmetric distribution of the 56Ni has also been proposed for these deviations from asphericity (Chugai, 2006; Dessart & Hillier, 2011). In the case of SN 2004dj, (Chugai et al., 2005) argued that the bipolar ejection of 56Ni led to the asymmetric excitation of hydrogen within a spherical hydrogen envelope due to the shock propagation not being purely radial. In such a scenario, a rapidly but smoothly changing PA might be expected during the transition from the photospheric phase to the nebular phase, during which 56Ni is uncovered. This could explain the PA rotation observed in SN 2023ixf from 68 d up until 140 d. Such an ejection of 56Ni is hypothesized to stem directly from bipolar explosions in CCSNe resulting from mildly rotating RSG cores (Chugai et al., 2005).
All the facets concerning the geometry described above are compatible with a jet-driven explosion, where the jet breaks out from the inner core into the hydrogen envelope (Wang et al., 2001). This would lead to an aspherical shock structure, not due to the interaction with the dense equatorial CSM but the asymmetric ejection of 56Ni. CCSNe are often associated with jets that puncture the outer layers or get choked by the hydrogen envelope (Couch et al., 2011). There are a few SNe II, like SN 2012aw (Kumar et al., 2014), SN 2013ej (Nagao et al., 2021) and SN 2017gmr (Nagao et al., 2019), which show an early onset of rise in polarization before the end of the plateau phase during the photospheric phase which is indicative of asymmetry in the hydrogen envelope, probably indicating a choked jet in the hydrogen envelope. The onset of polarization rise in SNe II appears to be linked to their luminosity, ejecta velocity, and 56Ni mass. CCSNe with higher luminosity, higher ejecta velocity, and increased 56Ni production exhibit more pronounced and prolonged aspherical characteristics (Nagao et al., 2024). This suggests that a higher likelihood of asphericity, possibly resembling a bipolar explosion leading to a jet-like formation, could be a crucial factor in the explosion mechanism of highly energetic SNe like SN 2023ixf.
An intriguing observation is that SN 2023ixf shows two preferred dominant axes: a) along the CSM-outer ejecta, where the SN evolved during the early phase until the middle of the plateau phase, and b) along the inner ejecta/He-core, where the SN evolved during the late-plateau and early nebular phase. In other words, we see a change in PA of 45 degrees from Peak B to Peak C in SN 2023ixf, suggesting that the irregularities in the explosion structure are not aligned with the orientation of the confined CSM. Therefore, if a jet-driven explosion drove SN 2023ixf, the rotation of the progenitor did not drive the jet. A similar change in PA was noticed in the polarization of SN 1987A (see Figure 11) from the photospheric phase to the early nebular phase, indicating a potential avenue for further investigation if hydrogen-rich SNe could be powered by jets. The -band imaging polarization evolution and spectropolarimetric observations of SN 2023ixf from close to the explosion (1.4 d) up until the early nebular phase (140 d) suggest significant departures from global sphericity in various facets of the evolution of SN 2023ixf, which encompasses the confined CSM, SN ejecta (possible the extended CSM), and the inner core.
9 Summary
SNe II are the most common type of CCSNe and yet harbor several mysteries regarding the late-stage evolution of its massive star progenitors, resulting in considerable observational heterogeneity. We present a comprehensive investigation of multi-wavelength photometry, optical spectroscopy, polarimetry, and spectropolarimetric observations of SN 2023ixf until the early nebular phase. We highlight the major results below:
Photometric evolution: SN 2023ixf showed rise times of 4.5 d (blue peak) arising due to CSM interaction and 16 d (red peak) arising from the SN ejecta. SN 2023ixf shows a early plateau decline rate (s1) of and a late-plateau decline rate (s2) of , resembling fast-declining SNe II. The plateau length of SN 2023ixf is 75 d, towards the shorter end of SNe II. SN 2023ixf is one of the brightest SN IIP/L ever observed in UV with a peak UVW1 magnitude of –20 mag.
Slow photospheric evolution and Distinctive profile: We infer a delayed development of metal features in the spectral sequence of SN 2023ixf, hinting at the ejecta cooling slower than a normal SNe II, possibly due to CSM interaction. The weaker absorption in the P-Cygni profile of suggests ongoing interaction during the plateau phase, reminiscent of Type IIL SNe. Post 16 d, the and are characterized by a high-velocity broad absorption feature at 13,500 in addition to the clumpy P-Cygni profile with an absorption minimum at 8,500 .
Confined CSM: The early discovery and classification allowed for the coverage of the SN just after 1 d wherein we observe an initial rise in the blackbody temperature evolution, blueward rise of UV colors, and order-of-magnitude rise in UV flux at a nearly constant radius of evolution asserting the delayed shock breakout due to a confined dense CSM in SN 2023ixf.
Hydrodynamical light curve modeling: Using STELLA to perform hydrodynamical light curve modeling of SN 2023ixf, we estimated a progenitor mass of 10 with a radius of 470 and an explosion energy of and 0.06 of . The inferred properties are not a unique solution since degeneracies exist in the modeling. The early UV excess was best modeled by a confined dense CSM spanning from the tip of the progenitor to arising from a mass-loss rate of 10-2 and a density structure. The late-plateau UV excess was modeled by an extended CSM spanning to with a mass-loss rate of 10-4 and a density structure. A wind-velocity of 10 was assumed for computing the mass-loss rates. The presence of the two-zone CSM in SN 2023ixf with 2 orders of magnitude difference in mass-loss rate suggests an increased mass-loss rate of the progenitor as it neared its explosion.
Asphericity in CSM, ejecta, and the He-core: The temporal evolution of ISP-corrected -band polarization in SN 2023ixf showed three distinct peaks: Point A [, ] at 1.4 d, Peak B [ at 6.4 d, ] and Peak C [, ] at 79.2 d. The observations indicate the presence of an asymmetric confined dense CSM, an aspherical shock front, and clumpiness in the low-density extended CSM, along with an aspherical inner He-core in SN 2023ixf. The inference on an aspherical shock front also stems from the dual broad absorption features seen in the blue wing of the Balmer features starting 16 d. Overall, SN 2023ixf displayed two dominant axes, one along the CSM-outer ejecta and the other along the inner ejecta/He-core, showcasing the independent origin of asymmetry during the early and late evolution.
Signs of Molecular CO / Dust formation: The flattening in the -band light curve and the attenuation of the red-edge of H post 125 d indicates early onset of molecular CO and dust formation in SN 2023ixf similar to SN 2017eaw and SN 1987A. However, our dataset lacks an NIR spectrum during the nebular phase to confirm the presence of molecular CO.
SN 2023ixf, the third closest CCSNe observed in the 21st century, merits deeper exploration to comprehensively analyze its high-cadence observations across the electromagnetic spectrum. Conducting late-phase nebular spectroscopy holds promise in illuminating the intricate, asymmetric structure of the CSM surrounding SN 2023ixf and providing insights into the properties of its synthesized dust. We anticipate that the aspherical shock structure may manifest in the observations at radio wavelengths, revealing evidence of both freely expanding and decelerated shocks. The interaction of the faster shock with the low-density CSM is expected to generate a rapid rise in the radio light curve, while the interaction with the equatorial dense CSM may result in a slower increase over the subsequent months. These distinct radio signatures prompt further observational follow-up to validate the findings outlined in this article.
10 Software and third party data repository citations
acknowledgments
We thank the anonymous referee for their comments which helped in the readability of the manuscript. We thank the IAO, Hanle, CREST, and Hosakote staff, who made the observations from HCT possible. The facilities at IAO and CREST are operated by the Indian Institute of Astrophysics, Bangalore. The GROWTH India Telescope (GIT) is a 70-cm telescope with a 0.7-degree field of view, set up by the Indian Institute of Astrophysics (IIA) and the Indian Institute of Technology Bombay (IITB) with funding from Indo-US Science and Technology Forum and the Science and Engineering Research Board, Department of Science and Technology, Government of India. It is located at the Indian Astronomical Observatory (IAO, Hanle). We acknowledge funding by the IITB alumni batch of 1994, which partially supports the operation of the telescope. Telescope technical details are available at GROWTH-India Website101010https://sites.google.com/view/growthindia/. This research is based partly on data obtained through the KASTOR (Kanata And Seimei Transient Observation Regime) program. The Seimei telescope at the Okayama Observatory is jointly operated by Kyoto University and the National Astronomical Observatory of Japan (NAOJ), with assistance provided by the Optical and Infrared Synergetic Telescopes for Education and Research (OISTER) program funded by the MEXT of Japan. The authors thank the TriCCS developer team (supported by the JSPS KAKENHI grant Nos. JP18H05223, JP20H00174, and JP20H04736 and by NAOJ Joint Development Research).
TJM is supported by the Grants-in-Aid for Scientific Research of the Japan Society for the Promotion of Science (JP24K00682, JP24H01824, JP21H04997, JP24H00002, JP24H00027, JP24K00668) and the Australian Research Council (ARC) through the ARC’s Discovery Projects funding scheme (project DP240101786). KM acknowledges support from the JSPS KAKENHI grant JP20H00174 and JP24H01810. DKS acknowledges the support provided by DST-JSPS under grant number DST/INT/JSPS/P 363/2022. GCA thanks the Indian National Science Academy for support under the INSA Senior Scientist Programme. BK acknowledges support from the “Science & Technology Champion Project” (202005AB160002) and from two “Team Projects” – the “Top Team” (202305AT350002) and the “Innovation Team” (202105AE160021), all funded by the “Yunnan Revitalization Talent Support Program”.
Numerical computations were in part performed by the PC cluster at the Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan. This research made use of RedPipe111111https://github.com/sPaMFouR/RedPipe (Singh, 2021), an assemblage of data reduction and analysis scripts written by AS. This research also made use of the NASA/IPAC Extragalactic Database (NED121212https://ned.ipac.caltech.edu), which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
The work used Swift/UVOT data reduced by P. J. Brown and released in the Swift Optical/Ultraviolet Supernova Archive (SOUSA). SOUSA is supported by NASA’s Astrophysics Data Analysis Program through grant NNX13AF35G. This work has also used software and/or web tools obtained from NASA’s High Energy Astrophysics Science Archive Research Center (HEASARC), a service of the Goddard Space Flight Center and the Smithsonian Astrophysical Observatory. This work was also partially supported by a Leverhulme Trust Research Project Grant.